OpenStudy (anonymous):

3*4^x-24=0. Please find x and do it step by step.

OpenStudy (anonymous):

3*4^x=24 (+24) 4^x = 8 (/3) 4^x= 4^1.5 x=1.5

jhonyy9 (jhonyy9):

4*x = 8 2*(2x) =2*(3) 2x=3 x=3/2

OpenStudy (anonymous):

\[3*4^x-24=0\] \[3*4^x=24\] \[4^x=8\] \[4^x=4^(3/2)\] \[x=3/2\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

3*4=81 x-21 81-21=60